OpenStudy (anonymous):

please help! solve for x; (1/2) + (4/2x) = (x+4/10)

OpenStudy (anonymous):

\[\frac{1}{2}+\frac{4}{2x}=\frac{x+4}{10}\]

OpenStudy (anonymous):

I would begin by using the law that \[\frac{a}{b}+\frac{c}{d}=\frac{ad + bc}{bd}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/2+2x=(x+4)/10 5+20x=x+4, 1=-19x, x=-1/19